OpenStudy (anonymous):

Find the rate of change between the two points: (8, 1) and (4, 2) where x is in months and y is in inches. 25 inches per month -.25 inches per month 4 inches per month -4 inches per month

OpenStudy (anonymous):

i think its -4 but im not sure

OpenStudy (anonymous):

\[\frac { 2-1}{4-8}= -\frac 1 4=-.25 \]

jimthompson5910 (jim_thompson5910):

rate of change = (change in y)/(change in x) change in y = 2 - 1 = 1 change in x = 4 - 8 = -4 rate of change = (change in y)/(change in x) rate of change = 1/(-4) rate of change = -1/4 rate of change = -0.25 So the rate of change is -0.25 inches per month

OpenStudy (anonymous):

is it neagtive or positive

jimthompson5910 (jim_thompson5910):

It's negative. It might help to plot the two points and draw a line through them. You'll see that this line slopes downward (as you read it from left to right).